# SAFE-SQL: Self-Augmented In-Context Learning with Fine-grained Example Selection for Text-to-SQL

### v1

@@ Abstract
Text-to-SQL aims to convert natural language questions into executable SQL queries. While previous approaches, such as skeleton-masked selection, have demonstrated strong performance by retrieving similar training examples to guide large language models (LLMs), they struggle in real-world scenarios where such examples are unavailable. To overcome this limitation, we propose Self-Augmentation in-context learning with Fine-grained Example selection for Text-to-SQL (SAFE-SQL), a novel framework that improves SQL generation by generating and filtering self-augmented examples. SAFE-SQL first prompts an LLM to generate multiple Text-to-SQL examples relevant to the test input. Then SAFE-SQL filters these examples through three relevance assessments, constructing high-quality in-context learning examples. Using self-generated examples, SAFE-SQL surpasses the previous zero-shot, and few-shot Text-to-SQL frameworks, achieving higher execution accuracy. Notably, our approach provides additional performance gains in extra hard and unseen scenarios, where conventional methods often fail.
@@
Jimin Lee,  Ingeol Baek,  Byeongjeong Kim,  Hwanhee Leethanks: Corresponding author. Department of Artificial Intelligence, Chung-Ang University, Seoul, Korea {ljm1690, ingeolbaek, michael97k, hwanheelee}@cau.ac.kr
@@ 1 Introduction
Text-to-SQL generation converts questions into SQL queries that help users access information in databases. Traditional approaches on Text-to-SQL, such as, rule-based systems and early machine learning models, relied on hand-crafted rules or simple pattern matching to generate SQL queries. They often struggle with the ambiguity and context-dependence of natural language, making it challenging to accurately translate user intent into structured SQL commands El Boujddaini et al. (2024); Mohammadjafari et al. (2025); Li and Jagadish (2014). As the field progressed, more sophisticated approaches emerged, including skeleton-masked selection Gao et al. (2023), relying on retrieving similar examples from training data to guide query generation.
Figure 1: The left example illustrates a failure in retrieving similar Text-to-SQL examples from the training set. In contrast, the right example demonstrates how our proposed self-augmented approach successfully generates similar examples autonomously.
@@
However, these methods face significant challenges in real-world scenarios where similar examples are often unavailable in the training set Gan et al. (2021); Hong et al. (2024) or retrieve unrelated examples as shown in Figure 1. To overcome these problems, recent research has introduced to generate synthetic data. SQL-GEN, introduced by Pourreza et al. (2024) introduces dialect-specific synthetic data to resolve the diverse SQL dialect challenges in Text-to-SQL systems. Another important aspect of synthetic data generation is incorporating key relationships from the schema and employing schema-distance-weighted column sampling Zhao et al. (2022). However, these synthetic data generation methodologies predominantly require supervised fine-tuning, which demands substantial computational resources and time Yang et al. (2024b). Moreover, self-generated examples can introduce significant noise and inaccuracies that undermine the quality of in-context learning. Errors in synthetic SQL queries or flawed reasoning paths may lead to incorrect interpretations of database schemas Wretblad et al. (2024). As a result, relying on unfiltered self-generated examples for Text-to-SQL tasks can risk degrading overall model performance. Consequently, it is necessary to develop more efficient approaches that enhance the accuracy of Text-to-SQL while eliminating extra training costs and mitigating the adverse impacts of noisy self-generated examples.
@@
In this paper, we propose SAFE-SQL, a novel approach that fully exploits the generative power of large language models (LLMs) to create high-quality synthetic examples in an unsupervised manner. SAFE-SQL enhances its inference capabilities without additional fine-tuning through four key steps: (1) Schema Linking: Analyzing SQL test questions, database tables, and foreign keys to map relationships between queries and database structures (2) Example Generation: Generating ten similar question-SQL query-reasoning path triplets per input using schema-linked information with LLMs (3) Threshold-based example selection: Evaluating generated examples via embedding similarity, keyword & structural alignment, and reasoning path validity, retaining only those scoring above specific threshold and (4) Final SQL Inference: Leveraging the curated examples, this step utilizes in-context learning to enhance the performance of large language models. This approach benefits from carefully selected examples that align with the natural language question and database schema, ensuring accurate and efficient SQL generation.
@@
By relying on LLM-generated and filtered examples, SAFE-SQL significantly improves robustness and accuracy, particularly in complex or unseen scenarios where retrieval-based approaches struggle. Consequently, our approach eliminates the need for additional model training while achieving superior performance in Text-to-SQL tasks. Our contributions can be listed as follows:
@@
  • We propose SAFE-SQL, a fully unsupervised approach that leverages LLMs to generate synthetic examples without additional fine-tuning or reliance on predefined training sets.
@@
  • We introduce a structured filtering mechanism that selects high-quality question-SQL pairs based on embedding similarity, keyword and structural alignment, and reasoning path validation.
  • Our method dynamically adapts examples using schema-linked information to boost SQL generation particularly in complex and unseen scenarios.
@@ 2 Related work
2.1 Supervised Fine-Tuning for Text-to-SQL
Supervised Fine-Tuning enables models to learn domain-specific nuances and schema alignments, significantly improving performance on specialized tasks Sarker et al. (2024). Techniques such as schema linking Yang et al. (2024c), constraint-based decoding Scholak et al. (2021), and execution-guided generation Wang et al. (2024) have further enhanced the robustness of fine-tuned models in handling domain-specific challenges. Synthesizes text-to-SQL data from weak and strong LLMs Yang et al. (2024b) utilizes preference learning from the weak data from small LLMs and strong data from Large LLMs. However, supervised fine-tuning can be time-consuming and demands substantial computational resources. To remedy such issues, our approach sidesteps these heavy requirements by generating high-quality synthetic examples in a fully unsupervised manner.
2.2 In-Context Learning for Text-to-SQL
In-Context learning has emerged as another influential method, leveraging the ability of large language models to perform text to sql tasks by conditioning on a few examples provided in the input prompt, without requiring explicit parameter updates. Unlike traditional in-context learning methods that rely on a fixed set of examples, our approach dynamically generates and filters schema-aware examples to better adapt to varying query structures in real-world scenarios. AST-SQL Shen et al. (2024) introduces using an abstract syntax tree algorithm to select similar examples and incorporate them into the in-context learning pipeline.
@@ 2.2 In-Context Learning for Text-to-SQL
Skeleton Masked Similarity Method
Skeleton masked similarity is an approach that emphasizes structural similarity between natural language questions and SQL queries. This method involves extracting the skeleton of a SQL query from the given question and masking unnecessary details to focus on its essential structure. By preserving key structural patterns, such as SELECT-FROM-WHERE clauses, this approach facilitates learning the correspondence between natural language expressions and SQL query elements. By moving beyond simple word-level matching, it captures deeper structural correlations, which is particularly effective for complex queries or linguistically diverse questions.
Classification and Decomposition Methods
The classification and decomposition method simplifies SQL query generation by breaking down natural language questions into sequential steps. This process starts with classifying the input based on its structure or intent (e.g., single-table queries, multi-table joins, or nested queries) and then decomposes the question into smaller subtasks, such as identifying specific clauses and resolving their components. PTD-SQL Luo et al. (2024) uses this step-wise approach to enhance LLM inference, reducing complexity and improving overall accuracy by modularizing the SQL generation process. While this method effectively handles complex and multi-intent queries, it struggles with questions that do not fit predefined categories.
Self-Training and Correction Methods
Recent advancements in the Text-to-SQL domain have focused on enhancing model performance and reliability through self-training and self-correction techniques. These approaches enable models to generate data or correct errors autonomously, thereby improving data efficiency and generalization capabilities. LG AI research introduces Self-training strategy Jo et al. (2024) utilizing pseudo-labeled unanswerable questions to develop a reliable Text-to-SQL system for Electronic Health Records (EHRs). SelECT-SQL Shen and Kejriwal (2024), proposes an innovative in-context learning solution that combines chain-of-thought prompting, self-correction, and ensemble methods. MAGIC Askari et al. (2024), use multiple agents (manager, feedback, and correction agents) to iteratively refine generated SQL queries. Other approaches employ code language models to repeatedly correct errors in SQL queries Chen et al. (2023). Unlike these iterative correction techniques, our approach preemptively filters and refines self-generated examples to reduce errors before query generation, thereby avoiding the need for additional correction steps.
@@ 3 Fine-grained Self-Augmentation for Text-to-SQL
We propose SAFE-SQL, a framework that automatically generates high-quality examples for in-context learning in Text-to-SQL tasks. Unlike traditional methods that rely on retrieving similar questions or using predefined templates, SAFE-SQL uses LLMs to create synthetic examples tailored to the given database schema. These examples are then filtered based on their semantic similarity, structural similarity, and the quality of their reasoning paths. Finally, we predict final SQL query for the test input using the self-generated examples via in-context learning.
@@ 3.1 Schema Linking
The first step in SAFE-SQL is schema linking, which identifies and extracts relevant schema elements from the database to reduce noise and improve performance in Text-to-SQL tasks Cao et al. (2024). As shown in Figure 2, the schema linking step involves analyzing the test question to detect keywords and phrases that correspond to schema elements such as tables, columns, rows, and foreign keys within the database schema. This mapping narrows the focus to the most pertinent parts of the schema and provides the necessary context for generating examples that are both meaningful and grounded in the database structure.
@@ 3.2 Example Generation
Using the information obtained from schema linking, the LLM generates multiple synthetic examples for each test question. As illustrated in Figure 2, for each test question, we generate ten examples—each comprising a similar question, its corresponding SQL query, and a detailed reasoning path. These examples maintain structural similarity while varying elements such as numerical values, table names, and key attributes. This controlled variation ensures that the generated examples remain relevant while encouraging the model to generalize beyond surface-level patterns. By observing these modified instances, the model can infer the correct SQL query even when faced with unseen but structurally similar questions. In particular, the reasoning path outlines the logical steps required to derive the correct SQL query result, providing a comprehensive explanation of the query execution process. We provide the full prompt used for Large Language Models in Appendix B.1.
@@ 3.3 Relevance Scoring
After generating synthetic examples, SAFE-SQL evaluates examples, as shown in Figure 2. This evaluation process capture the core intent of the test question. To achieve this, each example $e$ is assigned a composite relevance score $Rel$ on a scale from 0 to 10, which is calculated as follows:
@@
Here, $Q_{t}$ denotes the test question, $Q_{e}$ denotes the generated example question, $R$ denotes the reasoning path of the generated example, and the three components are defined as:
  • $S(Q_{e},Q_{t})$: Semantic similarity between the generated question and test questions.
  • $A(Q_{e},Q_{t})$: Structural alignment between the generated question and test questions.
  • $R$: Evaluates the quality of the reasoning path accompanying the example. This score measures the completeness, logical consistency, and alignment of the reasoning steps with both the test question and the relevant schema elements.
The weighting factors $\alpha$, $\beta$, and $\gamma$ sum to 1, allowing SAFE-SQL to select examples. Only examples that surpass a predefined threshold are retained for guiding SQL query generation. We provide the full prompt used for LLMs in Appendix B.2. The semantic similarity score measures whether the generated question accurately preserves the intent of the test question, even if its wording differs. As shown in Figure 2, a generated question "Retrieve the list of countries where one singer is older than 40 and another is younger than 30." has a different structure but correctly maintains the meaning, resulting in a high semantic similarity score. Another generated question, "Show cities where a singer above age 40 and a singer below 30 are from." shares the same SQL structure as the test question, differing only in the geographical entity, resulting in a high structural similarity score. For reasoning path similarity, examples that follow a similar step-by-step logical process to derive the SQL query receive a higher score. Queries that require the same table joins, filtering conditions, and aggregation methods are more likely to align with the reasoning process of the test question, leading to a stronger match. These factors are complementary, and carefully evaluating these factors is crucial for effective example selection, ensuring that the retrieved examples are highly relevant to the test question.
@@ 3.5 Final Inference
In the final stage, the high-quality examples generated in previous steps are combined with the test question to construct a comprehensive prompt for the LLM. These examples, enriched with similar questions, corresponding SQL queries, and detailed reasoning paths, guide the LLM in generating the final SQL query. By integrating schema linking, synthetic example generation, relevance scoring, and threshold-based filtering, SAFE-SQL produces SQL queries that are both syntactically correct and semantically aligned with the intended database operations, while also providing an interpretable reasoning process.
@@ 4.1 Experimental Setup
For our experiments, we employ six models for comparison purposes: GPT-4o Hurst et al. (2024), GPT-4o-mini Hurst et al. (2024), GPT-4 Achiam et al. (2023), Llama-3.1-8B-Instruct Dubey et al. (2024), Deepseek-coder-6.7b-instruct Guo et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024a), and starcoder2-7b Lozhkov et al. (2024). The evaluation is conducted on the Spider Yu et al. (2018) dev dataset, which is a widely used benchmark for Text-to-SQL systems. The Spider dev set is a large-scale, cross-domain benchmark specifically designed to assess the generalization capabilities of Text-to-SQL models. It contains 7,000 training samples covering 166 databases in various domains and 1,034 evaluation samples from 20 databases. Our analysis parts conduct with gpt-4o-2024-08-06 version model.
@@ 4.2 Baselines
We use the following baseline text to SQL methods: Supervised fine tuning, which fine-tune open source model, Zero-shot, which inference without examples, Few-shot, which inference with few examples. Synthesizes text-to-SQL data from weak and strong LLMs Yang et al. (2024b) utilizes preference learning from the weak data from small LLMs and strong data from LLMs. SQL-Palm Sun et al. (2024) introduces synthetic data augmentation to fine-tune open source models. Din SQL Pourreza and Rafiei (2023) breaking down the task into smaller sub-tasks, allowing large language models to iteratively improve their reasoning process through self-correction. C3 SQL Dong et al. (2023) comprises Clear Prompting, Calibration with Hints, and Consistent Output, which systematically addresses model input, bias, and output to enhance performance using zero-shot prompt. Dail SQL Gao et al. (2023) introduces effective few-shot learning which significantly reduces the number of tokens required per question. ACT-SQL Zhang et al. (2023) enhances Text-to-SQL performance by automatically generating chain-of-thought exemplars, eliminating the need for manual labeling. PTD SQL Luo et al. (2024) categorizing queries into subproblems and focusing on targeted drilling to improve large language models’ reasoning capabilities. Unlike traditional few-shot methods that depend on human-selected examples, our Self-Augmented In-Context Learning method allowing the model to generate its own in-context examples, and used for final inference. This self-augmented approach removing the need for manually provided exemples while still leveraging the benefits of in-context learning, leading to improved adaptability and performance in Text-to-SQL tasks.
Method Model Easy Medium Hard Extra All
@@
SYN-SQL Sense 13B 95.2 88.6 75.9 60.3 83.5
SQL-Palm PaLM2 93.5 84.8 62.6 48.2 77.3
@@
Baseline GPT-4 84.3 73.1 65.8 40.3 69.1
Baseline GPT-4o 87.2 77.2 68.4 48.7 73.4
Baseline GPT-4o-mini 84.8 75.6 67.0 46.1 71.5
C3q-SQL GPT-4 90.2 82.8 77.3 64.3 80.6
@@
SAFE-SQL GPT-4 93.2 88.9 85.8 74.7 86.8
SAFE-SQL GPT-4o 93.4 89.3 88.4 75.8 87.9
SAFE-SQL GPT-4o-mini 93.6 87.5 86.1 75.2 87.4

### v2

@@ Abstract
Text-to-SQL aims to convert natural language questions into executable SQL queries. While previous approaches, such as skeleton-masked selection, have demonstrated strong performance by retrieving similar training examples to guide large language models (LLMs), they struggle in real-world scenarios where such examples are unavailable. To overcome this limitation, we propose Self-Augmentation in-context learning with Fine-grained Example selection for Text-to-SQL (SAFE-SQL), a novel unsupervised framework that enhances SQL generation by generating and intelligently filtering self-augmented examples. SAFE-SQL leverages an LLM to generate diverse Text-to-SQL examples, which are then filtered by a novel fine-grained mechanism using criteria for semantic similarity, structural alignment, and reasoning path quality to curate high-quality in-context learning examples. Leveraging these carefully selected self-generated examples, SAFE-SQL significantly surpasses previous zero-shot and few-shot Text-to-SQL frameworks, achieving superior execution accuracy. Notably, our approach demonstrates substantial performance gains in challenging extra hard and unseen scenarios, where conventional methods often struggle.
@@
Jimin Lee,  Ingeol Baek,  Byeongjeong Kim  Hyunkyung Bae,  Hwanhee Leethanks: Corresponding author. Department of Artificial Intelligence, Chung-Ang University, Seoul, Korea {ljm1690, ingeolbaek, michael97k, hwanheelee}@cau.ac.kr
@@ 1 Introduction
Text-to-SQL generation converts questions into SQL queries that help users access information in databases. Traditional approaches on Text-to-SQL rely on hand-crafted rules or simple pattern matching to generate SQL queries. They often struggle with the ambiguity and context-dependence of natural language, making it challenging to accurately translate user intent into structured SQL commands El Boujddaini et al. (2024); Mohammadjafari et al. (2025); Li and Jagadish (2014). As the field progressed, more sophisticated approaches emerged, including skeleton-masked selection Gao et al. (2023), relying on retrieving similar examples from training data to guide query generation.
Figure 1: The example on the left shows a failure in retrieving relevant examples due to masked keywords, which results in superficially similar but actually unrelated questions being selected. In contrast, our self-augmented approach generates N-examples and filters them using 3 criteria, resulting in appropriate example retrieval.
@@
However, these methods face significant challenges in real-world scenarios where similar examples are often unavailable in the training set Gan et al. (2021); Hong et al. (2024) or unrelated examples are retrieved as shown in Figure 1. To overcome these problems, recent research has introduced methods to generate synthetic data. SQL-GEN, presents by Pourreza et al. (2024), introduces dialect-specific synthetic data to resolve the diverse SQL dialect challenges in Text-to-SQL systems. Another important aspect of synthetic data generation is incorporating key relationships from the schema and employing schema-distance-weighted column sampling Zhao et al. (2022). However, these synthetic data generation methodologies predominantly require supervised fine-tuning, which demands substantial computational resources and time Yang et al. (2024b). Moreover, self-generated examples can introduce significant noise and inaccuracies that undermine the quality of in-context learning. Errors in synthetic SQL queries or flawed reasoning paths may lead to incorrect interpretations of database schemas Wretblad et al. (2024). As a result, relying on unfiltered self-generated examples for Text-to-SQL tasks can pose a risk of degrading overall model performance. Consequently, it is necessary to develop more efficient approaches that enhance the accuracy of Text-to-SQL while eliminating extra training costs and mitigating the adverse impacts of noisy self-generated examples by implementing a robust filtering mechanism.
@@
In this paper, we propose SAFE-SQL, a novel approach that fully exploits the generative power of large language models (LLMs) to create high-quality synthetic examples in an unsupervised manner. SAFE-SQL enhances its inference capabilities without additional fine-tuning through four key steps: (1) Schema Linking: Analyzing SQL test questions, database tables, and foreign keys to map relationships between queries and database structures (2) Example Generation: Generating N-question-SQL query-reasoning path triplets per input using schema-linked information with LLMs (3) Threshold-based example selection: Filtering generated examples using specifically designed relevance criteria based on semantic similarity, Structural alignment, and reasoning path validity, retaining only those scoring above a specific threshold to ensure high quality and relevance for in-context learning examples and (4) Final SQL Inference: Leveraging the curated examples, this step utilizes in-context learning to enhance the performance of large language models. This approach benefits from carefully selected examples that align with the natural language question and database schema, ensuring accurate and efficient SQL generation.
@@
By relying on LLM-generated and filtered examples, SAFE-SQL significantly improves robustness and accuracy, particularly in complex or unseen scenarios where retrieval-based approaches struggle. Our approach eliminates the need for additional model training while achieving superior performance in Text-to-SQL tasks. Our contributions can be listed as follows:
@@
  • We propose SAFE-SQL, a fully unsupervised approach that leverages LLMs to generate synthetic examples.
  • Our method leverages schema linking to dynamically adapt examples, boosting the performance of Text-to-SQL in complex scenarios.
@@
  • We introduce a structured filtering mechanism that selects high-quality question-SQL pairs based on semantic similarity, structural alignment, and reasoning path validation.
@@ 2 Related Work
Structural and Semantic Information for Text-to-SQL
Advances in Text-to-SQL have increasingly emphasized the importance of effectively utilizing structural and semantic information derived from the database schema. RAT-SQL Wang et al. (2021) introduces relation-aware transformer architectures capable of encoding both the natural language question and the complex structure of the database schema, leading to improved schema linking performance. Concurrently, PICARD Scholak et al. (2021) demonstrates that leveraging constrained decoding with step-by-step execution during generation can reduce the likelihood of producing invalid SQL queries. Rather than treating SQL generation as a pure sequence prediction task, PICARD executes partial SQL statements during generation, thus enforcing syntactic and semantic correctness. Both of these approaches highlight the importance of integrating structural and semantic information to generate correct SQL queries. Building on this line of research, our work explores how self-augmented examples can effectively incorporate structurally and semantically relevant information for in-context learning, particularly in scenarios where large annotated datasets are unavailable.
In-context Learning with Example Augmentation and Filtering
As LLMs have demonstrated strong performance in in-context learning settings, recent work has focused on improving the effectiveness of demonstrations through better example augmentation and selection Toteja et al. (2025). Self-Instruct Wang et al. (2023) introduces a framework for generating instruction-tuning data by prompting the model to synthesize and filter examples, leveraging synthetic supervision. Further studies on demonstration selection for in-context learning Wang et al. (2024) have systematically studied strategies for selecting effective in-context examples, including similarity-based retrieval and clustering methods. Integrating self-generated prompts with explicit reasoning chains has also been shown to significantly improve in-context learning outcomes by guiding the model’s thought process Shum et al. (2023); Wei et al. (2023). These studies collectively underscore the impact of demonstration quality on LLM performance and highlight the potential of intelligently curating examples. Focusing on the Text-to-SQL task, our work distinguishes itself by generating synthetic examples and filtering them using a novel fine-grained mechanism that considers semantic similarity, structural similarity, and reasoning path quality. This allows us to generate multiple candidate examples and select the most effective through this tailored fine-grained filtering process.
@@ 3 Fine-grained Self-Augmentation for Text-to-SQL
We propose SAFE-SQL, a framework that automatically generates high-quality examples for in-context learning in Text-to-SQL tasks. Unlike traditional methods that rely on retrieving similar questions or using predefined templates, SAFE-SQL uses LLMs to create synthetic examples tailored to the given database schema. These examples are then filtered based on their semantic similarity, structural alignment, and the quality of reasoning paths. Finally, we predict the final SQL query for the test input using the self-generated examples via in-context learning.
@@ 3.1 Schema Linking
The first step in SAFE-SQL is schema linking, which identifies and extracts relevant schema elements from the database to reduce noise and improve performance in Text-to-SQL tasks Cao et al. (2024). As shown in Figure 2, the schema linking step involves analyzing the test question to detect keywords and phrases that correspond to schema elements such as tables, columns, rows, and foreign keys within the database schema. This mapping narrows the focus to the most pertinent parts of the schema and provides the necessary context for generating relevant examples that are both meaningful and grounded in the database structure.
@@ 3.2 Example Generation
Using the information obtained from schema linking, the LLM generates a pool of multiple synthetic examples for each test question. As illustrated in Figure 2, for each test question, we generate ten examples—each comprising a similar question, its corresponding SQL query, and a detailed reasoning path. The generated SQL questions maintain structural similarity while varying elements such as numerical values, table names, and key attributes. This ensures that the generated examples remain relevant while encouraging the model to generalize beyond surface-level patterns. By observing these modified instances, the model can infer the correct SQL query even when face with unseen but structurally similar questions. In particular, the reasoning path outlines the logical steps required to derive the correct SQL query result, providing a comprehensive explanation of the query execution process. We provide the full prompt used for LLMs in Appendix B.1.
@@ 3.3 Relevance Scoring
After generating a set of synthetic examples, SAFE-SQL employs a crucial evaluation process rooted in novel fine-grained example selection to determine the relevance of each generated example to the test question. This fine-grained selection process is integral to our method, ensuring that only high-quality, contextually appropriate examples are used for in-context learning, moving beyond simple retrieval to curate examples truly relevant and beneficial for the Text-to-SQL task.
To achieve this, we assign composite relevance score $Rel$ on a scale from 0 to 10 to each example $e$, which is calculated as follows:
@@
Here, $Q_{t}$ represents the test question, $Q_{e}$ denotes the generated example question. The coefficients $\alpha$, $\beta$, and $\gamma$ are weighting factors that sum to 1, allowing for adjustment of the relative importance of each component in the fine-grained selection score. The three components are defined as follows:
  • Semantic Similarity $S(Q_{e},Q_{t})$: assesses if the generated question preserves the underlying meaning and intent to ensure the example aligns with the user’s core query objective.
  • Structural Alignment $A(Q_{e},Q_{t})$: evaluates structural correspondence based on key database elements and their relationships, which is important for mapping natural language to a similar database structure and operations.
  • Reasoning Path Quality $R$: evaluates the alignment of the example’s logical derivation steps and database operations (e.g., filtering, aggregation, joins, subqueries) with the test question’s required logic.
We utilize LLMs to compute the score for each of the three components based on our specifically designed instructions and criteria. Specifically, the LLM applies a predefined, multi-point scoring rubric (detailed in Appendix B.2) to assign a quantitative score (0-10) for each criterion. This process allows for a nuanced assessment of the degree of alignment between the generated example and the test question along each dimension, moving beyond simple binary or qualitative judgments. By carefully evaluating these three factors through our fine-grained example selection process, SAFE-SQL ensures that the selected examples are highly relevant and informative, contributing to more accurate and effective SQL query generation.
@@ 3.5 Final Inference
In the final stage, the high-quality examples generated in previous steps are combined with the test question to construct a comprehensive prompt for the LLM. These examples, enriched with filtered questions, corresponding SQL queries, and detailed reasoning paths, guide the LLM in generating the final SQL query. By integrating schema linking, synthetic example generation, relevance scoring, and threshold-based filtering, SAFE-SQL produces SQL queries that are both syntactically correct and semantically aligned with the intended database operations, while also providing an interpretable reasoning process.
@@ 4.1 Experimental Setup
For our experiments, we employ six models for comparison purposes: GPT-4o Hurst et al. (2024), GPT-4o-mini Hurst et al. (2024), GPT-4 Achiam et al. (2023), Llama-3.1-70B-Instruct Dubey et al. (2024), Llama3.3-70B-Instruct Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024a), Gemma3-12b Team et al. (2025), and Gemma3-27b Team et al. (2025). The evaluation is conducted on the Spider dev dataset Yu et al. (2018) and the Bird dev dataset Li et al. (2023), which are widely used benchmarks for Text-to-SQL systems. The Spider dev set contains 7,000 training samples covering 166 databases in various domains and 1,034 evaluation samples from 20 databases, comprised of four difficulty levels. BIRD is a large cross-domain Text-to-SQL dataset with 12,751 question-SQL pairs across 95 databases. Since the test sets of both the Spider and BIRD datasets are only accessible through specific evaluation servers, we conduct our evaluation using their respective development sets.
@@ 4.2 Baselines
We use the following baseline Text-to-SQL methods: Supervised fine tuning, which fine-tunes an open source model, Zero-shot, which infers without examples, Few-shot, which infers with few examples. Synthesizing Text-to-SQL data from weak and strong LLMs Yang et al. (2024b) utilizes preference learning from the weak data from small LLMs and strong data from LLMs. SQL-Palm Sun et al. (2024) introduces synthetic data augmentation to fine-tune open source models. Din SQL Pourreza and Rafiei (2023) breaking down the task into smaller sub-tasks, allowing large language models to improve their reasoning process through self-correction iteratively. C3-SQL Dong et al. (2023) comprises clear prompting, calibration with hints, and consistent output, which systematically addresses model input, bias, and output to enhance performance using the zero-shot prompt. Dail-SQL Gao et al. (2023) introduces effective few-shot learning, significantly reducing the number of tokens required per question. ACT-SQL Zhang et al. (2023) enhances Text-to-SQL performance by automatically generating chain-of-thought exemplars, eliminating the need for manual labeling. PTD-SQL Luo et al. (2024) categorizes queries into subproblems and focuses on targeted drilling to improve LLMs’ reasoning capabilities.
Method Model Easy Medium Hard Extra All Time
@@
SYN-SQL Sense 13B 95.2 88.6 75.9 60.3 83.5 16.34
SQL-Palm PaLM2 93.5 84.8 62.6 48.2 77.3 18.93
@@
Baseline GPT-4 84.3 73.1 65.8 40.3 69.1 1.28
Baseline GPT-4o 87.2 77.2 68.4 48.7 73.4 0.93
Baseline GPT-4o-mini 84.8 75.6 67.0 46.1 71.5 1.07
C3-SQL GPT-4 90.2 82.8 77.3 64.3 80.6 19.34
@@
SAFE-SQL GPT-4 93.2 88.9 85.8 74.7 86.8 21.41
SAFE-SQL GPT-4o 93.4 89.3 88.4 75.8 87.9 14.92
SAFE-SQL GPT-4o-mini 93.6 87.5 86.1 75.2 87.4 15.33
SAFE-SQL Llama3.1-70B-Instruct 90.4 88.2 86.2 78.2 86.8 23.52
SAFE-SQL Llama3.3-70B-Instruct 92.0 80.5 81.0 62.9 80.5 22.46
SAFE-SQL Qwen2.5-72B 87.6 74.5 77.0 52.4 74.5 28.51
SAFE-SQL Gemma3-12B 92.4 90.7 85.1 78.8 88.2 13.28
SAFE-SQL Gemma3-27B 93.6 89.8 87.4 78.8 88.5 14.26